# Efficient Decoder for End-to-End Oriented Object Detection in Remote Sensing Images

### v1

@@ III-C Selective Distinct Queries
RRoIFormer R50 87.45 78.57 47.36 69.01 79.58 81.27 88.53 90.89 82.80 86.21 58.68 64.20 75.21 74.44 61.39 75.04
RRoIFormer* R50 87.33 81.90 54.54 72.65 80.82 83.70 88.97 90.82 83.48 85.57 62.27 59.00 79.07 81.88 66.39 77.23
IV EXPERIMENTS
IV-A Datasets
We conduct our experiments on five public datasets, including oriented object detection datasets DIOR-R [38], DOTA-v1.0 [8]/v1.5/v2.0 [49], and a horizontal object detection dataset MS COCO [9].
DIOR-R [38] is a large-scale oriented object detection dataset. It contains a total of 23463 remote sensing images and 192512 instances covering 20 common classes including windmill (WM), trainstation (TS), storagetank (STO), ship (SH), harbor (HA), golffield (GF), expressway toll station (ETS), bridge (BR), baseballfield (BF), airplane (APL), vehicle (VE), tenniscourt (TC), stadium (STA), overpass (OP), groundtrackfield (GTF), dam (DAM), expressway service area (ESA), chimney (CH), basketballcourt (BC) and airport (APO). We train our model on training and validation sets and report the results on the test set.
DOTA series including DOTA-v1.0/v1.5/v2.0 are used. DOTA-v1.0 [8] is designed for oriented object detection tasks which contain 188282 instances and 2806 remote sensing images ranging in size from 800×800 to 4000×4000. It covers 15 different object categories: Plane (PL), Baseball diamond (BD), Bridge (BR), Ground track field (GTF), Small vehicle (SV), Large vehicle (LV), Ship (SH), Tennis court (TC), Basketball court (BC), Storage tank (ST), Soccer-ball field (SBF), Roundabout (RA), Harbor (HA), Swimming pool (SP), and Helicopter (HC). DOTA-v1.5 uses the same images as DOTA-v1.0 but adds a new category Container Crane (CC) and more small instances, which contains 403318 instances. DOTA-v2.0 adds two categories Airport (Air) and Helipad (Heli) compared to DOTA-v1.5, which contains 11268 images and 1793658 instances. We use both training and validation sets for training and test set for testing for the DOTA series. The detection result of the test set is obtained by submitting testing results to DOTA’s official evaluation server.
MS COCO 2017 [9] is a horizontal object detection dataset for natural scenes, consisting of 118K images as train2017 for training and 5K images as val2017 for validation. We train our model on train2017 and report the evaluation results on val2017.
IV-B Implementation Details
We select ResNet-50 [28] as the backbone which is pre-trained on the ImageNet [50]. We optimize the overall network with AdamW optimizer [51] with the warming up for 500 iterations. We use Focal loss [35] for classification and Rotated IoU/GIoU loss [36, 37] and L1 loss for regression.
For experiments on the DOTA-v1.5/v2.0, we use 2 NVIDIA RTX 2080ti with a batch size of 4 (2 images per GPU) and the learning rate of $5\times e^{-5}$ for training, while a single GPU with a batch size of 2 and the learning rate of $2.5\times e^{-5}$ on DOTA-v1.0 and DIOR-R. Models build on MMRotate [52] with Pytorch [53]. The weights of losses are 2.0, 2.0 and 5.0 for Focal loss, L1 loss and Rotated IoU loss respectively. Following official settings of the DOTA benchmark, we crop images into patches of $1024\times 1024$ with overlaps of 200 and train the model for 24 epochs. Images on DIOR-R are trained for 36 epochs with the original fixed size of $800\times 800$, while 12 epochs in ablation studies. To compare with other methods, the number of queries is 500, while in ablation studies on DIOR-R, it is 300.
For experiments on COCO 2017, we use 2 NVIDIA RTX 2080ti with a batch size of 4 (2 images per GPU) for training. Models build on MMDetection [54] with Pytorch. The weights of loss are 2.0, 5.0, and 2.0 for Focal loss, L1 loss, and IoU loss respectively. The learning rate is $2\times e^{-5}$.
IV-C Main Results
Results on DIOR-R. As shown in Table II, our proposed RRoIFormer is compared with CNN-based one-stage and two-stage detectors and transformer-based detectors. It achieves the state-of-the-art performance of 67.31% mAP with ResNet50 backbone on the DIOR-R benchmark. The backbones of all models are ResNet50 in Table II. Our method surpasses all comparison CNN-based one-stage and two-stage detectors and Transformer-based detectors. Overall, our method obtains 0.51%, 2.9%, and 1.41% improvements in mAP over DCFL [1], AOPG [38], and ARS-DETR [42].
Results on DOTA-v1.0. We compare our RRoIFormer method with modern CNN-based one-stage and two-stage detectors and transformer-based detectors. Table III reports the detailed results of every category on the DOTA-v1.0 dataset. With ResNet50 as the backbone, RRoIFormer obtains 75.04% mAP under single-scale training and testing. In addition, with the multi-scale training augmentation strategy, our method achieves 77.23% mAP using the ResNet50 backbone. RRoIFormer surpasses most comparison methods, particularly compared to Transformer-based detectors.
Results on DOTA-v1.5. We report results on DOTA-v1.5 in Table IV compared with current CNN-based one-stage and two-stage detectors. The backbones of all models are ResNet50 in Table IV. Our method achieves 67.43% mAP which achieves state-of-the-art performance under single-scale training and testing. Especially for the categories with small instances (e.g., HA, SP, CC), it performs better.
Results on DOTA-v2.0. Tabel V shows a comparison of our method with the modern detectors on the DOTA-v2.0 dataset. The backbones of all models are ResNet50 in Table V. Our proposed method achieves the new state-of-the-art performance of 53.28% mAP under single-scale training and testing. It is comparable to the mAP of the previous best models Oriented R-CNN [3].
TABLE IV: Comparison with state-of-the-art methods on the DOTA-v1.5 dataset.
Mehtod PL BD BR GTF SV LV SH TC BC ST SBF RA HA SP HC CC mAP
one-stage:
RetinaNet-O [35] 71.43 77.64 42.12 64.65 44.53 56.79 73.31 90.84 76.02 59.96 46.95 69.24 59.65 64.52 48.06 0.83 59.16
DCFL [1] - - - - 56.72 - 80.87 - - 75.65 - - - - - - 67.37
two-stage:
Faster RCNN-O [27] 71.89 74.47 44.45 59.87 51.28 68.98 79.37 90.78 77.38 67.50 47.75 69.72 61.22 65.28 60.47 1.54 62.00
Mask R-CNN [55] 76.84 73.51 49.90 57.80 51.31 71.34 79.75 90.46 74.21 66.07 46.21 70.61 63.07 64.46 57.81 9.42 62.67
HTC [56] 77.80 73.67 51.40 63.99 51.54 73.31 80.31 90.48 75.12 67.34 48.51 70.63 64.84 64.48 55.87 5.15 63.40
ReDet [4] 79.20 82.81 51.92 71.41 52.38 75.73 80.92 90.83 75.81 68.64 49.29 72.03 73.36 70.55 63.33 11.53 66.86
end-to-end:
RRoIFormer 78.44 75.43 48.84 63.62 62.16 78.41 88.94 90.84 75.29 81.83 57.05 66.44 73.95 72.69 51.64 13.37 67.43
TABLE V: Comparison with state-of-the-art methods on the DOTA-v2.0 dataset.
Method Plane BD Bridge GTF SV LV Ship TC BC ST SBF RA Harbor SP HC CC Air Heli mAP
one-stage:
SASM [45] 70.30 40.62 37.01 59.03 40.21 45.46 44.60 78.58 49.34 60.73 29.89 46.57 42.95 48.31 28.13 1.82 76.37 0.74 44.53
Oriented Rep [39] 73.02 46.68 42.37 63.05 47.06 50.28 58.64 78.84 57.12 66.77 35.21 50.76 48.77 51.62 34.23 6.17 64.66 5.87 48.95
ATSS-O [26] 77.46 49.55 42.12 62.61 45.15 48.40 51.70 78.43 59.33 62.65 39.18 52.43 42.92 53.98 42.70 5.91 67.09 10.68 49.57
DCFL [1] 75.71 49.40 44.69 63.23 46.48 51.55 55.50 79.30 59.96 65.39 41.86 54.42 47.03 55.72 50.49 11.75 69.01 7.75 51.57
two-stage:
Mask R-CNN [55] 76.20 49.91 41.61 60.00 41.08 50.77 56.24 78.01 55.85 57.48 36.62 51.67 47.39 55.79 59.06 3.64 60.26 8.95 49.47
HTC [56] 77.69 47.25 41.15 60.71 41.77 52.79 58.87 78.74 55.22 58.49 38.57 52.48 49.58 56.18 54.09 4.20 66.38 11.92 50.34
RoI Transformer [34] 71.81 48.39 45.88 64.02 42.09 54.39 59.92 82.70 63.29 58.71 41.04 52.82 53.32 56.18 57.94 25.71 63.72 8.70 52.81
Oriented R-CNN [3] 77.95 50.29 46.73 65.24 42.61 54.56 60.02 79.08 61.69 59.42 42.26 56.89 51.11 56.16 59.33 25.81 60.67 9.17 53.28
end-to-end:
RRoIFormer 77.58 49.96 38.60 53.82 54.97 57.12 68.93 77.88 59.59 71.92 40.17 51.22 53.00 57.12 49.66 25.42 66.92 5.17 53.28
TABLE VI: Results for SDQ applied to Sparse R-CNN.
Model w/SDQ #query #epochs mAP FPS Params(M)
Sparse R-CNN 100 12 37.9 20.6 106
100 12 41.0 (+3.1) 20.0 106
300 36 45.0 20.5 106
300 36 46.8 (+1.8) 19.3 106
TABLE VII: Ablation study of the effects of proposed invidual stategy.
Method RRoI attn SDQ mAP Params(M) FPS
baseline 61.66 64 13.4
RRoIFormer 61.67 41 13.9
62.76 41 14.1
IV-D Effects of Selective Distinct Queries
To verify the effectiveness of our proposed Selective Distinct Queries module, we apply the SDQ to a horizontal object detector Sparse R-CNN [13] with the ResNet50 backbone. The model is trained and tested on COCO 2017 and detailed results are reported in table VI. We collect selective queries from only $5^{th}$ decoder layer with a total of 6 decoder layers. Then we obtain distinct queries also in front of the last decoder layers and the IoU threshold is set to 0.7. With 100 queries and 12 training epochs, SDQ significantly improves +3.1% mAP compared with the baseline. Increasing the number of queries to 300 and training epoch to 36, SDQ achieves 46.8% (+1.8%) mAP on the COCO val2017 set with the help of a random resize augmentation strategy from DETR [6]. These experiments can fully prove the effectiveness of the SDQ method. The SDQ would not bring extra parameters of the model and FPS is only slightly reduced (20.0 vs 20.6).
IV-E Ablation Study
@@
Effects of Individual Strategy. We check the effectiveness of each proposed strategy in the proposed method. As shown in table VII, the baseline detector DDQ-O [12] yields a result of 61.66% mAP on DIOR-R with 300 queries. When we gradually apply the RRoI attention and SDQ into the baseline detector, our model achieves 61.67% and 62.76% mAP respectively, verifying the effectiveness of each design. RRoI attention significantly reduces the parameters by 23 million compared to dynamic interaction in the baseline detectors. When applying RRoI attention to the baseline model, due to the reduction in parameters, the FPS increases by 0.5. When we further apply SDQ, since there are fewer distinct queries than the initial queries, the FPS increases slightly by 0.2.
The number of attention head. Following deformable attention [7], our RRoI attention also adopts multiple heads. Different heads establish different associations between queries and values. In this ablation, we use different numbers of $m$ heads. As we increase the number of heads from 2 to 8, the mAP grows from 61.36% to 62.63% on DIOR-R, shown in Table VIII. This indicates that the multi-head attention extends the ability to focus on different parts of features, and provides multiple subspaces for representation. When we increase the number of heads to 16, there is a decline in performance. We suspect that the redundant head focuses on redundant features.
Different pooling size. To further explore RRoI Attention, we adopt different pooling sizes $r$. As the pooling size $r$ increases, more features will be highlighted in oriented boxes. When pooling size $r$ is set to 5, the AP is only 61.35% on DIOR-R, shown in Table IX. When it is increased to 7, the mAP can achieve 62.63%. This suggests that abundant features are sufficient to represent objects. When we increase the pooling size to 9, the performance of model declines. It indicates that redundant features confuse the attention mechanism.
TABLE VIII: Different number of attention heads of RRoI Attention.
Head 2 4 8 16
mAP 61.36 62.19 62.63 55.63
TABLE IX: Different pooling size of RRoI Attention.
Pooling Size 5 7 9
mAP 61.35 62.63 62.32
TABLE X: Ablation study of different IoU threshold for SDQ.
threshold 0.8 0.85 0.9 0.95
mAP 62.58 62.76 62.63 62.49
TABLE XI: Ablation study of different decoder layer in Selective Distinct Queries.
Method Layers Selective Distinct AP
SDQ 3 $0^{th}$, $1^{st}$, $2^{nd}$ $2^{nd}$ 58.01
$1^{st}$, $2^{nd}$ $2^{nd}$ 60.66
$1^{st}$ $1^{st}$ 58.03
$2^{nd}$ $2^{nd}$ 62.08
TABLE XII: Ablation study of reducing the decoder layers.
Method Layers Selective Distinct AP
SDQ 3 $2^{nd}$ $2^{nd}$ 62.08
2 $1^{st}$ $1^{st}$ 62.63
1 $0^{th}$ $0^{th}$ 44.02
TABLE XIII: Ablation study of different query selection mechanism.
Method Layers Selective Distinct AP
Top-k 3 $1^{st}$, $2^{nd}$ $2^{nd}$ 59.49
Stack 59.99
SDQ 60.66
Threshold of Selective Distinct Queries. To further explore the effectiveness of SDQ, we conducted ablation experiments on the different IoU thresholds shown in table X. To filter the similar queries and keep the remaining queries distinct, the threshold is a high value close to 1 in SDQ. As the IoU threshold value increases from 0.8 to 0.85, the mAP increases by 0.18% on DIOR-R. When the IoU threshold increases to 0.95 which is very close to 1, the mAP value decreases slightly, indicating that there are very few predicted boxes that fully coincide or are close to perfectly coinciding.
Different decoder layers in Selective Distinct Queries. To further verify the effectiveness of SDQ, we performed ablation experiments with different numbers of decoder layers in SDQ on DIOR-R, shown in table XI. The total number of decoder layers is 3 here. When we collect queries from the $2^{nd}$ decoder layer as selective queries and filter similar queries to obtain distinct queries also at the $2^{nd}$ layer, the model achieves the best 62.08% mAP on DIOR-R. When selective queries contain more queries from front layers, e.g. initial queries, there is a significant decline in performance. Queries at the front layers can not adequately represent objects.
Reduce total decoder layers. To further verify the effectiveness of SDQ, we performed ablation experiments on reducing the total decoder layers on DIOR-R, shown in Table XII. When only one decoder layer, we do SDQ on initial queries. In this condition, our method only achieves 44.02% mAP. When we increase the number of decoder layers from 2 to 3, the mAP drops from 62.63% to 62.08%. The decoder and attention bring too many parameters which may be redundant.
Queries selection mechanism. To further verify the effectiveness of SDQ, we performed ablation experiments with different selection mechanisms on DIOR-R, including top-$k$ and stack mechanisms, shown in table XIII. Following DINO [10], we choose top-$k$ queries according to scores of multi-class classification, while in Deformable DETR [7], the top-$k$ method selects proposals only according to scores of binary classification. In the top-$k$ mechanism, we collect queries from $1^{st}$ and $2^{nd}$ decoder layers and choose top-$k$ queries at $2^{nd}$ layer. The number of $k$ is equal to initial queries. In addition, we densely stack queries from $1^{st}$ and $2^{nd}$ decoder layers to $3^{th}$ layers which can be regarded as a revision of SQR [22]. The number of queries multiplies many times according to candidate layers through the stack. Essentially, the method of stack increases performance by increasing the number of queries. For a fair comparison, our SDQ also collects selective queries from $1^{st}$ and $2^{nd}$ decoder layers. Compared with the other two different methods, our SDQ mechanism is the most effective and is the only one that can reduce the number of queries.

### v4

@@ 4.3 Ablation Study
RQFormer R50 87.45 78.57 47.36 69.01 79.58 81.27 88.53 90.89 82.80 86.21 58.68 64.20 75.21 74.44 61.39 75.04
RQFormer* R50 87.98 86.13 57.79 82.12 81.33 84.87 88.78 90.90 88.65 87.45 70.74 70.39 80.33 81.03 75.31 80.92
SDQ vs. other queries selection mechanism. To further verify the effectiveness of SDQ, we performed ablation experiments with different selection mechanisms on DIOR-R, including top-$n$ and stack mechanisms, shown in Table 2 (e). Following DINO (Zhang et al., 2023a), we choose top-$n$ queries according to scores of multi-class classification, while in Deformable DETR (Zhu et al., 2020), the top-$n$ method selects proposals only according to scores of binary classification. In the top-$n$ mechanism, we collect queries from $1^{st}$ and $2^{nd}$ decoder layers and choose top-$n$ queries at $2^{nd}$ layer. The number of $n$ is equal to initial queries. In addition, we densely stack queries from $1^{st}$ and $2^{nd}$ decoder layers to $3^{th}$ layers which can be regarded as a revision of SQR (Chen et al., 2023a). The number of queries multiplies many times according to candidate layers through the stack. Essentially, the method of stack increases performance by increasing the number of queries. For a fair comparison, our SDQ also collects selective queries from $1^{st}$ and $2^{nd}$ decoder layers. Compared with the other two different methods, our SDQ mechanism is the most effective and is the only one that can reduce the number of queries.
Table 5: Main results of small size objects on DOTA-v1.5. The results in bold denote the best performance of each column.
Method SV LV SH ST SP CC mAP
RetinaNet-O 44.53 56.79 73.31 59.96 64.52 0.83 59.16
DCFL 56.72 - 80.87 75.65 - - 67.37
Faster RCNN-O 51.28 68.98 79.37 67.50 65.28 1.54 62.00
Mask R-CNN 51.31 71.34 79.75 66.07 64.46 9.42 62.67
HTC 51.54 73.31 80.31 67.34 64.48 5.15 63.40
ReDet 52.38 75.73 80.92 68.64 70.55 11.53 66.86
RQFormer 62.16 78.41 88.94 81.83 72.69 13.37 67.43
Table 6: Performance comparisons on the DOTA-v2.0 dataset.
Method SASM RetinaNet-O Oriented Rep Mask R-CNN
mAP 44.53 46.68 48.95 49.47
Method ATSS-O S2A-Net HTC DCFL
mAP 49.57 49.86 50.34 51.57
Method RoI Trans. S2A-Net$+$DCFL Oriented R-CNN RQFormer
mAP 52.81 52.84 53.28 53.28
Table 7: Comparison of the performance of different methods on ICDAR2015. ’P’ is precision and ’R’ is recall. The results in bold denote the best performance of each column. The backbone used by all methods is Resnet50.
Method P R F-measure FPS
R3DET 71.7 79.8 75.5 -
GWD 74.0 80.5 77.1 -
PSC 84.2 73.9 78.7 20.4
Oriented Reppoints 84.3 74.0 78.8 21.4
CFA 85.3 73.5 78.9 21.4
S2A-Net 80.4 78.2 79.3 -
Oriented RCNN 82.9 77.3 80.0 10.0
DDQ-O (baseline) 85.1 73.8 79.0 14.1
RQFormer 85.0 75.5 80.0 14.1
Number of decoder layers. Multiple decoder layers are a widely used technique to improve detection performance. We performed ablation experiments on different decoder layers on DIOR-R, to further verify the effectiveness of SDQ, shown in Table 2 (f). When only one decoder layer, we do SDQ on initial queries. In this condition, our method only achieves 44.02% mAP. When we increase the number of decoder layers from 2 to 3, the mAP drops from 62.63% to 62.08%. The decoder and attention bring too many parameters which may be redundant.
Different combinations of selective queries and distinct queries in SDQ. To further verify the effectiveness of SDQ, we performed ablation experiments with different combinations of selective queries and distinct queries in SDQ on DIOR-R, shown in Table 2 (g). The total number of decoder layers is 3 here. When we collect queries from the $2^{nd}$ decoder layer as selective queries and filter similar queries to obtain distinct queries also at the $2^{nd}$ layer, the model achieves the best 62.08% mAP on DIOR-R. When selective queries contain more queries from front layers, e.g., initial queries, there is a significant decline in performance. Queries at the front layers can not adequately represent objects.
@@
Effects of Individual Strategy. We check the effectiveness of each proposed module in the proposed method. In Table 2 (h), the baseline detector DDQ-O (Zhang et al., 2023b) produces an mAP of 61.66% on DIOR-R with 300 queries. Sequentially incorporating RRoI attention and SDQ into the baseline detector results in mAP of 62.31% and 62.41%, respectively, affirming the effectiveness of each design. When both modules are applied, the model achieves an mAP of 62.76%. Notably, RRoI attention significantly reduces parameters by 23 million compared to instance interaction in the baseline. Introducing RRoI attention to the baseline model, due to the parameter reduction, increases the FPS by 0.36. Subsequent application of SDQ, requiring the filtering of similar queries, leads to a slight decrease in FPS. Importantly, SDQ brings performance improvement without introducing additional parameters.
4.4 Main Results
Results on DIOR-R. As shown in Table 3, our proposed RQFormer is compared with CNN-based one-stage and two-stage detectors and transformer-based detectors, including RetinaNet-O (Lin et al., 2017), DFDet (Xie et al., 2024), SOOD (Xi et al., 2024), Oriented Reppoints (Li et al., 2022b), DCFL (Xu et al., 2023), Gliding Vertex (Xu et al., 2020), RoI Transformer (Ding et al., 2019), QPDet (Yao et al., 2022), AOPG (Cheng et al., 2022), ARS-DETR (Zeng et al., 2024), OrientedFormer (Zhao et al., 2024), and DDQ-O (Zhang et al., 2023b). For fairness comparison, the backbones of all models are ResNet50 in Table 3. It achieves the promising performance of 67.31% mAP with ResNet50 backbone on the DIOR-R benchmark. Our method surpasses all comparison CNN-based one-stage and two-stage detectors and Transformer-based detectors. Specifically, our method outperforms DCFL (Xu et al., 2023) by 0.51% (67.31% vs. 66.80%), AOPG (Cheng et al., 2022) by 2.90% (67.31% vs. 64.41%), and ARS-DETR (Zeng et al., 2024) by 1.41% (67.31% vs. 65.90). We reconstruct DDQ-O (Zhang et al., 2023b) as our baseline and it achieves an mAP of 66.51%. Our RQFormer outperforms baseline by 0.80% (67.31% vs. 66.51%). Furthermore, RQFormer performs better in some categories with large aspect ratios, e.g., airport (APO) and dam. Take the category of dam for an example, RQFormer improves Oriented RepPoints (Li et al., 2022b), AOPG (Cheng et al., 2022), ARS-DETR (Zeng et al., 2024) and DDQ-O (Zhang et al., 2023b) with the gains of 5.28%, 8.77%, 0.95%, and 6.48%. It proves that RQFormer can detect accurately objects with extreme aspect ratios.
Table 8: Evaluation Results on MS COCO 2017. The proposed SDQ is applied to horizontal object detector Sparse R-CNN and DDQ.
Model w/SDQ query epochs mAP AP50 AP75 FPS Params
Sparse R-CNN 100 12 37.9 56.0 40.5 20.6 106M
100 12 41.0 ($\uparrow$ 3.1) 59.4 44.4 20.0 106M
300 36 45.0 64.1 48.9 20.5 106M
300 36 46.8 ($\uparrow$ 1.8) 65.9 51.2 19.3 106M
DDQ 100 12 42.1 59.8 46.3 9.5 64M
100 12 42.4 ($\uparrow$ 0.3) 59.6 46.5 9.4 64M
Table 9: Effectiveness of individual modules in RQFormer. DIOR-R and DOTA-v1.0 are used in this experiment.
Method RRoI Att. SDQ DIOR-R DOTA-v1.0
mAP Params(M) FPS mAP Params(M) FPS
baseline 66.51 63.73 14.02 73.57 63.71 11.27
RQFormer 67.11 40.93 14.42 74.05 40.90 11.82
RQFormer 67.27 63.73 13.54 74.89 63.71 11.08
RQFormer 67.31 40.93 14.00 75.04 40.90 11.20
Table 10: Speed, Parameters and accuracy on DOTA-v1.0.
Method Frame Backbone FPS Params (M) mAP
RoI Transformer Two- stage R50 9.18 55.35 74.61
Oriented RCNN R50 7.30 41.37 75.87
Gliding Vertex R101 10.23 41.36 75.02
R3Det One- stage R101 6.12 42.13 73.79
CFA R50 16.59 36.83 74.51
SASM R50 15.87 36.83 74.92
PSC R50 16.81 36.77 71.83
DCFL R50 17.73 36.10 74.26
D.DETR-O End- to- end R50 10.83 40.82 63.42
D.DETR-O w/CSL R50 10.71 41.15 72.15
ARS-DETR R50 10.12 41.56 73.79
DDQ-O (baseline) End- to-end R50 11.27 63.71 73.57
RQFormer R50 11.20 40.90 75.04
Results on DOTA-v1.0. We compare our RQFormer method with modern CNN-based one-stage and two-stage detectors and transformer-based detectors, including TIR-Net (Li et al., 2024), R3Det (Yang et al., 2021a), CFA (Guo et al., 2021), SASM (Hou et al., 2022), PSC (Yu and Da, 2023), H2RBox (Yang et al., 2022), SCRDet (Yang et al., 2019), RoI Transformer (Ding et al., 2019), Gliding Vertex (Xu et al., 2020), Oriented R-CNN (Xie et al., 2021), Deformable DETR-O (Zhu et al., 2020), EMO2-DETR (Hu et al., 2023) and ARS-DETR (Zeng et al., 2024). Table 4 reports the detailed results of every category on the DOTA-v1.0 dataset. With ResNet50 as the backbone, RQFormer obtains 75.04% mAP under single-scale training and testing. In addition, with the multi-scale training augmentation strategy, our method achieves 80.92% mAP using the ResNet50 backbone. RQFormer surpasses all CNN-based one-stage and transformer-based methods. Specifically, with single-scale data, RQFormer outperforms SASM (Hou et al., 2022) by 0.12% (75.04% vs. 74.92%) and ARS-DETR (Zeng et al., 2024) by 1.25% (75.04% vs. 73.79%). In addition, RQFormer outperforms most CNN-based two-stage methods, except for the Oriented R-CNN (Xie et al., 2021). We argue that it benefits from the complex architecture. Different from these two-stage models, our method does not require hand-craft postprocessing.
Results on DOTA-v1.5. We report results of small-size objects on DOTA-v1.5 in Table 5. We compare our method with RetinaNet-O (Lin et al., 2017), DCFL (Xu et al., 2023), Faster RCNN-O (Ren et al., 2015), Mask R-CNN (He et al., 2017), HTC (Chen et al., 2019), and ReDet (Han et al., 2021b). The backbones of all models are ResNet50. Our RQFormer performs better for the categories with small instances. Taking small vehicles as an example, our method improves DCFL (Xu et al., 2023) and ReDet (Han et al., 2021b) with the mAP gains of 5.44% and 9.78%, respectively. It proves the effectiveness of our method on small objects.
Results on DOTA-v2.0. Table 6 shows a comparison of our method with the modern detectors on the DOTA-v2.0 dataset. We compare our method with SASM (Hou et al., 2022), RetinaNet-O (Lin et al., 2017), Oriented Reppoints (Li et al., 2022b), Mask R-CNN (He et al., 2017), ATSS-O (Zhang et al., 2020), S2A-Net (Han et al., 2021a), HTC (Chen et al., 2019), DCFL (Xu et al., 2023), RoI Transformer (Ding et al., 2019), and Oriented R-CNN (Xie et al., 2021). The backbones of all models are ResNet50 in Table 6. Our proposed method achieves the competitive performance of 53.28% mAP under single-scale training and testing. It is comparable to the mAP of the previous best models Oriented R-CNN (Xie et al., 2021).
Results on ICDAR2015. Text detection involves locating text within images using boxes. It is worth noting that, as an extension of RQFormer, we only utilize rotated rectangular boxes to locate texts. We conduct experiments using the ICDAR2015 (Karatzas et al., 2015) dataset, shown in Table 7. We compare our method with R3DET (Yang et al., 2021a), GWD (Yang et al., 2021b), PSC (Yu and Da, 2023), Oriented Reppoints (Li et al., 2022b), CFA (Guo et al., 2021), S2A-Net (Han et al., 2021a), Oriented RCNN (Xie et al., 2021), and DDQ-O (Zhang et al., 2023b). The precision, recall, and F-measure are the evaluation metrics following official criteria (Karatzas et al., 2015). We reimplement compared methods using MMRotate (Zhou et al., 2022b), with the same settings as the RQFormer. The backbone used by all methods is ResNet50 (He et al., 2016). The proposed RQFormer achieves a precision of 85.0%, a recall of 75.5%, and an F-measure of 80.0%. It enhances the detection performance by 1.0% compared with the baseline and outperforms all the compared CNN-based models. The inference speed of RQFormer is 14.1 FPS, surpassing the two-stage method Oriented RCNN (10.0 FPS), but lagging behind the one-stage method CFA (21.4 FPS). Due to the limitation of rotating rectangular boxes, oriented object detection methods are not as flexible and accurate as dedicated text detection methods.
Table 11: Effect of positional and content queries on DIOR-R. Gradually increasing the number of queries, the performance is the best when it is 500.
Query 100 200 300 400 500 600 700
mAP 60.23 61.57 62.76 65.36 67.31 66.84 66.51
Params 40.8M 40.8M 40.9M 40.9M 40.9M 40.9M 41.0M
Figure 6: The ratios of similar queries under different high IoU threshold on DOTA-v1.0. Compared with others, our method significantly filters similar queries.
Figure 7: Convergence curves of Deformable DETR-O with CLS, ARS-DETR, DDQ-O and RQFormer with ResNet50 on DIOR-R.
4.5 Generalization to More Tasks
Results on COCO 2017. To verify the generalization of our method in horizontal object detection, we also conduct experiments on the MS COCO 2017 (Lin et al., 2014) dataset. Due to horizontal boxes, RRoI attention is not applicable in this task. Thus, we only apply the Selective Distinct Queries (SDQ) module to the horizontal object detectors Sparse R-CNN (Sun et al., 2021) and DDQ (Zhang et al., 2023b) with the ResNet50 backbone to verify the effectiveness of SDQ. The model is trained and tested on COCO 2017 and detailed results are reported in table 8. In Sparse R-CNN, we collect selective queries only from the $5^{th}$ decoder layer out of a total of 6 decoder layers, while in DDQ, we collect them from the $1^{st}$ decoder layer out of a total of 2 decoder layers. We then obtain distinct queries before the final decoder layers, with the IoU threshold set to 0.7. With 100 queries and 12 training epochs, SDQ significantly improves mAP by +3.1% compared to Sparse R-CNN and by +0.3% compared to DDQ. Increasing the number of queries to 300 and training epoch to 36, SDQ achieves 46.8% (+1.8%) mAP on the Sparse R-CNN (Sun et al., 2021) with the help of a random resize augmentation strategy (Carion et al., 2020). These experiments can fully prove the effectiveness of the SDQ method. The SDQ would not bring extra parameters of the model and FPS is only slightly reduced (20.0 vs. 20.6).
4.6 Effectiveness of individual modules
To further investigate the effectiveness of the individual proposed module, we gradually integrate RRoI attention and SDQ into the baseline DDQ-O (Zhang et al., 2023b), training them on both DIOR-R and DOTA-v1.0. Model evaluation involves mAP, parameters, and FPS, shown in Table 9. In this experiment, the number of queries is set to 500, and models are trained for 36 epochs on DIOR-R and 24 epochs on DOTA-v1.0. The FPS is tested on an RTX 2080ti. The baseline DDQ-O achieves a mAP of 66.51% and 73.57% on DIOR-R and DOTA-v1.0, respectively. When only incorporating RRoI attention, models exhibit an mAP improvement of 0.5% on DIOR-R and 0.48% on DOTA-v1.0. RRoI attention significantly reduces model parameters (40.93M vs. 63.73M on DIOR-R), as it focuses only on local areas within boxes. Fewer parameters result in a slight increase in FPS. With the integration of only SDQ, models demonstrate an mAP improvement of 0.76% on DIOR-R and 1.32% on DOTA-v1.0. SDQ does not introduce extra parameters to models, benefiting from its capability to filter similar queries. In this condition, the FPS experiences a slight decrease since SDQ first collects queries from intermediate layers and then filters similar queries. When both RRoI attention and SDQ are applied on the baseline, RQFormer achieves a mAP of 67.31% and 75.04% on DIOR-R and DOTA-v1.0, respectively.
4.7 Comparison of speed, parameters and accuracy
To further assess the performance of RQFormer, we conducted a comparison with other methods in terms of FPS, parameters, and accuracy on the DOTA-v1.0 dataset, shown in Table 10. RQFormer achieved a detection speed of 11.20 FPS, surpassing all compared CNN-based two-stage methods, but trailing behind most one-stage methods. We attribute the slower speed of CNN-based two-stage models to their complex structure and the influence of handcrafted post-processing modules. On the other hand, CNN-based one-stage models, characterized by a simpler framework and strategy than end-to-end models, tend to be faster. Notably, the speed of RQFormer is faster than all compared end-to-end methods, which we attribute to the elimination of an encoder and fewer decoder layers in our approach. The parameters of RQFormer amount to 40.90M, significantly less than baseline DDQ-O (Zhang et al., 2023b) (63.71M), while achieving a higher mAP. The parameters of RQFormer are also fewer than most compared CNN-based two-stage and end-to-end methods, though slightly larger than most one-stage methods. It proves the efficiency of our method in terms of parameters.
4.8 Effects of Queries
We performed ablation experiments with different number of queries, which including positional and content queries, as shown in Table 11. The number of positional queries and content queries is always equal. As the number of queries increases from 100 to 500, the mAP improves from 60.23% to 67.31%. Since remote sensing images contain a large number of densely packed objects, having a sufficient number of queries allows for better coverage of these objects. The number of parameters in the model only increases slightly, by 0.1M.